OpenStudy (anonymous):

Problem: Compute the total number Sn of different Morse code messages (message=any string of dots and dashes) that can be sent in n time units? To solve this problem, complete the following steps: 1.Show that S1=1, S2=2, S3=3, S4=5 by just listing all of the possible messages in each case.

OpenStudy (anonymous):

2.Explain why the following recurrence relation holds for Sn: Sn = Sn-1 + Sn-2 for n > 2 3.Find a formula for Sn as a function of n; that is, solve the discrete dynamical system: (I) S1 = 1, S2 = 2 (R) Sn = Sn-1 + Sn-2 for n > 2